OpenStudy (anonymous):

Which of the following equations have the same solution? Solve all three equations separately, and explain each step. I. 2x + 4 = 3x – 11 II. 3x + 3 = 51 III. the quantity of 2x plus 1, divided by 3 equals the quantity of x plus 6, divided by 2

OpenStudy (anonymous):

2x + 4 = 3x - 11 -- subtract 3x 2x - 3x + 4 = -11 -- subtract 4 2x - 3x = -11 - 4 -- combine like terms -x = - 15 -- multiply by -1 to make x positive x = 15 ******************* 3x + 3 = 51 -- subtract 3 3x = 51 - 3 3x = 48 -- divide by 3 x = 16 ******************* (2x + 1) / 3 = (x + 6) / 2 cross multiply 2(2x + 1) = 3(x + 6) -- distribute 4x + 2 = 3x + 18 -- subtract 3x 4x - 3x + 2 = 18 -- subtract 2 4x - 3x = 18 - 2 -- combine like terms x = 16 *******************

OpenStudy (anonymous):

thank you!! very much!! :)

OpenStudy (anonymous):

no problem :)